# Proper maps of annuli

### v1

@@ 2. Basic results on proper maps of annuli
Let us start with proving some of the basic statements we made in the introduction. First, we have the classical result about proper mappings of annuli in one dimension.
@@
Let us sketch the proof as it may not be easy to find, but it is almost identical to classifying biholomorphisms of annuli. First using an inversion and scaling and the maximum principle, we may assume that $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert$ goes to 1 as we approach the outer circle and it goes to $R$ as we approach the inner. The map $z\mapsto\frac{\log\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert}{\log R}$ is continuous on $\overline{\mathbb{A}_{1,r}}$, harmonic inside and goes to 0 on the outer circle and 1 on the inner circle. Hence it is equal to $\frac{\log\mathopen{}\mathclose{{\left\lvert z}}\right\rvert}{\log r}$, and we get $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{\frac{\log R}{\log r}}$. So $f(z)=e^{i\theta}z^{d}$ and $R=r^{d}$ for some $d\in\mathbb{N}$. ∎
In [helal-2025-proper], the authors have proved, using the theorem of Forstnerič [forstneric-1989-extending] as a key ingredient, that all annulus maps are rational when $n\geq 2$. For completeness we reproduce it here.
Theorem 2.2 ([helal-2025-proper]).
@@
Unlike in the ball case, no hypothesis on the boundary regularity is needed due to the pseudoconcavity of the inner boundary. The best expression for $D$ is not known, though D’Angelo has conjectured [dangelo-2003-sharp] that it is $\frac{N-1}{n-1}$ if $n\geq 3$ and 2N32N-3 is $n=2$.
@@
Via a theorem of Cima–Suffridge [cima-1990-boundary], a rational proper map of balls has no poles on the spheres, so we get a sphere map. Conversely a rational map taking a sphere to a sphere is a proper ball map via maximum principle. In [helal-2025-proper], it was proved that a rational sphere map also takes the outside of the sphere to the outside of the sphere. Putting all this together one gets the following equivalence.
@@
Next, let us show that a nonrational proper map exists from $\mathbb{A}_{1,r}$ to $\mathbb{A}_{2,R}$ for some $R$. First, we construct a nonrational proper map from the disc $\mathbb{D}$ to the 2-ball $\mathbb{B}_{2}$. Such a map follows from the result of Dor [dor-1990-proper], but it is useful to consider an elementary construction in the one-dimensional case. Take an arbitrary nonrational function from the disc $f_{1}\colon\mathbb{D}\to\mathbb{D}$, suppose that it extends continuously to the closed disc $\overline{\mathbb{D}}$ but does not extend holomorphically past any point on $\partial\mathbb{D}$ and suppose that $\sup_{z\in\overline{\mathbb{D}}}\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert<1$. Find a real-valued harmonic function $u$ on $\mathbb{D}$ continuous on $\overline{\mathbb{D}}$ so that on $\partial\mathbb{D}$ we have $u(z)=\frac{1}{2}\log\mathopen{}\mathclose{{\left(1-\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}}}\right)$. Find its harmonic conjugate $v\colon\mathbb{D}\to\mathbb{R}$. Then write $f_{2}=e^{u+iv}$. Then the mapping $f=f_{1}\oplus f_{2}$ is a proper map from $\mathbb{D}$ to $\mathbb{B}_{2}$ because $\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+\mathopen{}\mathclose{{\left\lvert f_{2}(z)}}\right\rvert^{2}=\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+e^{2u}\to 1$ as $\mathopen{}\mathclose{{\left\lvert z}}\right\rvert\to 1$. Clearly $f$ is not spherically equivalent to a rational map if $f_{1}$ is not rational.
@@
Let us compare the effect of the automorphism groups for proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ versus $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. The automorphism group of $\mathbb{B}_{n}$, $\operatorname{Aut}(\mathbb{B}_{n})$, is the set of linear fractional transformations of the form (9) below. For any affine subspace $L$ of dimension $m<n$ such that $L\cap\mathbb{B}_{n}\not$$\varnothing$, there exists a $\varphi\in\operatorname{Aut}(\mathbb{B}_{n})$ so that $\varphi(L)$ is given by $z_{m+1}=\cdots=z_{n}=0$.
@@
The effect of this difference on classification up to equivalence with respect to the embedding dimension $N_{f}$ is the following. For a proper ball map $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ where $N_{f}<N$, there exists an automorphism $\psi\in\operatorname{Aut}(\mathbb{B}_{N})$ so that $\psi\circ f=F\oplus 0$ where $F\colon\mathbb{B}_{n}\to\mathbb{B}_{N_{f}}$ is a proper map ($N_{f}=N_{F}$). The corresponding result for annulus maps is somewhat different, as an automorphism of the annulus is a unitary and cannot take an affine space to a linear space, it can only “rotate” it. The proof is an elementary computation.
@@
(8) $Uf=\sqrt{1-c^{2}}\,F\oplus c\oplus 0$
@@
In the proposition, the 0 in the direct sum has $N-N_{f}-1$ components (so it can be left out if $N=N_{f}+1$). The proposition leads to classification of first degree proper rational maps. We first start with the equidimensional case. In this case, unitaries (so unitarily equivalent to the identity) are the only maps.
@@
We have that $f$ is an LFT that must take the unit sphere to the unit sphere, hence it is well-known that it must have the form
(9) $U\phi_{a}(z)=U\frac{a-L_{a}z}{1-\langle z,a\rangle}\qquad L_{a}z=\mathopen{}\mathclose{{\left(1-\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}}}\right)\frac{\langle z,a\rangle}{\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}a+\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}z,$
where $U$ is a unitary and $a\in\mathbb{B}_{n}$. We need to show that $a=0$. By precomposing with a unitary, we can assume that $a=(a_{1},0,\ldots,0)$ and
(10) $\phi_{a}(z)=\mathopen{}\mathclose{{\left(\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}},\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{2}}{1-z_{1}\bar{a}_{1}},\ldots,\frac{-\sqrt{1-\mathopen{}\mathclose{{\left\lvert a_{1}}}\right\rvert^{2}}\,z_{n}}{1-z_{1}\bar{a}_{1}}}}\right).$
We note that the complex line $z_{2}=\cdots=z_{n}=0$ gets taken by $\phi_{a}$ to this same complex line, and on this line $\phi_{a}$ is simply the corresponding LFT in one dimension $\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$.
It remains to show that if $z_{1}\mapsto\frac{a_{1}-z_{1}}{1-z_{1}\bar{a}_{1}}$ takes a circle of radius $r<1$ centered at zero to a circle of radius $R<1$ centered at zero. In other words, we must have that for all $\theta$, $\mathopen{}\mathclose{{\left\lvert a_{1}-re^{i\theta}}}\right\rvert^{2}=R^{2}\mathopen{}\mathclose{{\left\lvert 1-re^{i\theta}\bar{a}_{1}}}\right\rvert^{2}$. By expanding and equating coefficients one obtains that $a_{1}=0$ and $r=R$. In other words, $a=0$ and $f$ is a unitary. ∎
As a quick corollary we can remove the hypothesis of rationality and degree. By Hartogs phenomenon and the maximum principle, we find that a holomorphic proper map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ extends to a proper holomorphic map of $\mathbb{B}_{n}$ to itself. That this map is a rational map of degree 1 then follows by Alexander’s theorem.
@@
When $N>n$, the result is a little more complicated.
@@
In the proof we will use the notion of a rank of a real polynomial $r(z,\bar{z})$ in $\mathbb{C}^{n}$. By this rank we mean the rank of its matrix of coefficients, or alternatively the smallest number $k$ so that
(12) $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2},$
where $\epsilon_{j}\pm 1$ and $P_{j}$ are holomorphic polynomials.
@@
Write $f=\frac{p}{q}$. As $f$ takes the unit sphere to the unit sphere we have
(13) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lVert q(z)}}\right\rVert^{2}=\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$
where $\lambda$ has to be a constant as $p$ and $q$ are of degree 1. Therefore, the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ has to be $n$ as that is the number of positive eigenvalues of the matrix of coefficients $\lambda(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1)$. In other words, $N_{f}=n$ and we can apply Proposition˜2.4. The $F$ in the proposition is unitarily equivalent to the identity by Proposition˜2.5. The values of the factors follow by elementary computation. ∎
We derive a few of basic results about the $k_{f}$ integer. First we note that for any rational map, it is enough to check an open dense subset of hyperplanes, that is, the general hyperplane. Similar but more complicated statement holds for holomorphic maps defined in some domain, but we have no need for the complication. The space of affine hyperplanes in $\mathbb{C}^{n}$ has the topology of the Grassmannian (minus the hyperplane at infinity).

### v2

@@ 2. Basic results on proper maps of annuli
Let us start with proving some of the basic statements we made in the introduction. First, we have the classical result about proper maps of annuli in one dimension.
@@
Let us sketch the proof as it may not be easy to find, but it is almost identical to classifying biholomorphisms of annuli. After a possible inversion and scaling, we assume that $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert$ goes to 1 as we approach the outer circle and it goes to $R$ as we approach the inner circle. The map $z\mapsto\frac{\log\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert}{\log R}$ is continuous on $\overline{\mathbb{A}_{1,r}}$, harmonic inside and is 0 on the outer circle and 1 on the inner circle. Hence it is equal to $\frac{\log\mathopen{}\mathclose{{\left\lvert z}}\right\rvert}{\log r}$, and we get $\mathopen{}\mathclose{{\left\lvert f(z)}}\right\rvert=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{\frac{\log R}{\log r}}$. So $f(z)=e^{i\theta}z^{d}$ and $R=r^{d}$ for some $d\in\mathbb{N}$. ∎
In [19], the authors have proved, using the theorem of Forstnerič [17] as a key ingredient, that all annulus maps are rational when $n\geq 2$. For completeness we reproduce it here.
Theorem 2.2 ([19]).
@@
Unlike in the ball case, no hypothesis on the boundary regularity is needed due to the pseudoconcavity of the inner boundary. The best expression for $D(n,N)$ is not known, though D’Angelo has conjectured [3] that for ball maps, it is $\frac{N-1}{n-1}$ if $n\geq 3$ and 2N32N-3 if $n=2$. For annulus maps, we know that the $D$ must be smaller. For example, by our result for $n=2$ and $N=3$, we see that $D(2,3)=2$ in the annulus case, while Faran shows that in the ball case the degree bound is $3$. We conjecture that for annulus maps in dimension $n\geq 2$ and degree $d$, we have $N\geq\binom{n+d-1}{d}$.
@@
Via a theorem of Cima–Suffridge [2], a rational proper map of balls has no poles on the spheres, so we get a sphere map. Conversely, a rational map taking a sphere to a sphere is a proper ball map via maximum principle. In [19], it was proved that a rational sphere map also takes the outside of the sphere to the outside of the sphere. Putting all these together one gets the following equivalence.
@@
Next, let us show that a nonrational proper map exists from $\mathbb{A}_{1,r}$ to $\mathbb{A}_{2,R}$ for some $R$. First, we construct a nonrational proper map from the disc $\mathbb{D}$ to the 2-ball $\mathbb{B}_{2}$. Such a map follows from the result of Dor [13], but it is useful to consider an elementary construction in the one-dimensional case. Take an arbitrary nonrational function from the disc $f_{1}\colon\mathbb{D}\to\mathbb{D}$, suppose that it extends continuously to the closed disc $\overline{\mathbb{D}}$ but does not extend holomorphically past any point on $\partial\mathbb{D}$ and suppose that $\sup_{z\in\overline{\mathbb{D}}}\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert<1$. Find a real-valued harmonic function $u$ on $\mathbb{D}$ continuous on $\overline{\mathbb{D}}$ so that on $\partial\mathbb{D}$ we have $u(z)=\frac{1}{2}\log\mathopen{}\mathclose{{\left(1-\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}}}\right)$. Find its harmonic conjugate $v\colon\mathbb{D}\to\mathbb{R}$ and write $f_{2}=e^{u+iv}$. Then the map $f=f_{1}\oplus f_{2}$ is a proper map from $\mathbb{D}$ to $\mathbb{B}_{2}$ because $\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+\mathopen{}\mathclose{{\left\lvert f_{2}(z)}}\right\rvert^{2}=\mathopen{}\mathclose{{\left\lvert f_{1}(z)}}\right\rvert^{2}+e^{2u}\to 1$ as $\mathopen{}\mathclose{{\left\lvert z}}\right\rvert\to 1$. If $f_{1}$ is not rational, then $f$ is not spherically equivalent to a rational map.
@@
Let us compare the effect of the automorphism groups for proper holomorphic maps $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ versus $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$. The automorphism group of $\mathbb{B}_{n}$, $\operatorname{Aut}(\mathbb{B}_{n})$, is the set of linear fractional transformations of the form
(8) $U\phi_{a}(z)=U\frac{a-L_{a}z}{1-\langle z,a\rangle}\qquad L_{a}z=\mathopen{}\mathclose{{\left(1-\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}}}\right)\frac{\langle z,a\rangle}{\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}a+\sqrt{1-\mathopen{}\mathclose{{\left\lVert a}}\right\rVert^{2}}z.$
For any affine subspace $L$ of dimension $m<n$ such that $L\cap\mathbb{B}_{n}\not$$\varnothing$, there exists a $\varphi\in\operatorname{Aut}(\mathbb{B}_{n})$ so that $\varphi(L)$ is given by $z_{m+1}=\cdots=z_{n}=0$.
@@
The effect of this difference on classification up to equivalence with respect to the embedding dimension $N_{f}$ is the following. For a proper ball map $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ where $N_{f}<N$, there exists an automorphism $\psi\in\operatorname{Aut}(\mathbb{B}_{N})$ so that $\psi\circ f=F\oplus 0$ where $F\colon\mathbb{B}_{n}\to\mathbb{B}_{N_{f}}$ is a proper map ($N_{f}=N_{F}$). The corresponding result for annulus maps is somewhat different, as an automorphism of the annulus is a unitary and cannot take an affine space to a linear space; it can only “rotate” it. The proof is an elementary computation.
@@
(9) $Uf=\sqrt{1-c^{2}}\,F\oplus c\oplus 0,$
@@
In the proposition, the 0 in the direct sum has $N-N_{f}-1$ components (so it can be left out if $N=N_{f}+1$).
To classify annulus maps the following division lemma is very useful. If $f=\frac{p}{q}$ is a rational proper annulus map, then it takes the unit sphere to the unit sphere and the $r$-sphere to the $R$-sphere. Namely, $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}$ is divisible by $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1$ and $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}-R^{2}\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}$ is divisible by $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}$. The following lemma encapsulates both of these conditions. The lemma is a special case of Lemma 3.11 from [19].
Lemma 2.5.
Suppose $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{N,R}$ is a proper rational map of degree $d$, $n\geq 2$, and write $f=\frac{p}{q}$ in lowest terms. Then $q$ is of degree at most d1d-1 and
(10) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(1+b\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\Bigr)\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
for a real polynomial $Q(z,\bar{z})$ and $b=\frac{1-R^{2}}{1-r^{2}}$. The $Q$ is of degree at most d2d-2 in $z$ and d2d-2 in $\bar{z}$.
Let us use the lemma to classify maps of degree 1. We consider the equidimensional case. In the proof we will use the notion of a rank of a real polynomial $\rho(z,\bar{z})$ in $\mathbb{C}^{n}$. By this rank we mean the rank of its matrix of coefficients, or alternatively the smallest number $k$ so that
(11) $\rho(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2},$
where $\epsilon_{j}=\pm 1$ and $P_{j}$ are holomorphic polynomials.
@@
We apply Lemma˜2.5 to find $b$ and $Q$. As $f$ is of degree 1, then $Q\equiv 0$ and $q$ is a constant, meaning we can just take $q=1$. The rank of $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$ must be the target dimension $n$, but the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=(1-b)+b\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$ is $n+1$ unless $b=1$. So $b=1$ and $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$, which implies that $f(z)=Uz$ for a unitary $U$. ∎
As a quick corollary, we can remove the hypothesis of rationality and degree. By Hartogs phenomenon and the maximum principle, we find that a proper holomorphic map $f\colon\mathbb{A}_{n,r}\to\mathbb{A}_{n,R}$ extends to a proper holomorphic map of $\mathbb{B}_{n}$ to itself. That this map is a rational map of degree 1 then follows by Alexander’s theorem.
@@
When $N>n$, the result is a little more complicated but follows in the same way.
@@
Again apply Lemma˜2.5 to find $q=1$ and $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=(1-b)+b\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}$. For this to be a squared norm we must have $b\leq 1$, which implies $R\geq r$. The form then follows as above. ∎
3. General hyperplane rank and its connection to the $X$-variety
Before proving some basic results about $k_{f}$ that we will require, let us discuss some of the history of this invariant. When Dor [13] constructed a nonrational map of balls, to show nonrationality, he showed that (in our language) $k_{f}=N$, as any rational ball map has $k_{f}\leq N-1$. When $k_{f}=N-1$, we obtain a map of Grassmannians that takes an open dense set of affine hyperplanes in $\mathbb{C}^{n}$ to affine hyperplanes in $\mathbb{C}^{N}$. This technique was used by Faran to classify proper maps from $\mathbb{B}_{2}$ to $\mathbb{B}_{3}$. The number $k_{f}$ is also related to the $X$-variety that Forstnerič used to show the rationality result. Let us quickly define the $X$-variety and note the connection for rational maps. Let $f\colon\mathbb{B}_{n}\to\mathbb{B}_{N}$ be a rational proper map. By complexifying the relation that $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}=1$ whenever $\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}=1$, we find $\langle f(z),f(w)\rangle=1$ whenever $\langle z,w\rangle=1$. The $X$-variety is defined as
(13) $X_{f}=\{(z,\zeta)\in\mathbb{C}^{n}\times\mathbb{C}^{N}:\langle\zeta,f(w)\rangle=1\text{ for all $w$ such that }\langle z,w\rangle=1\}.$
Note that the equation for (almost) every affine hyperplane in $\mathbb{C}^{n}$ can be written as $\langle z,w\rangle=1$ for a fixed $w$. Moreover, $\langle\zeta,f(w)\rangle=1$ gives affine hyperplanes on the target. The fiber of the $X$-variety is the subset for a fixed $z$. Such a fiber is an affine subspace of some dimension. It is not difficult to show that if the generic fiber of the $X$-variety has dimension $\ell$, then $k_{f}=N-\ell-1$. The $X$-variety encodes everything about the map $f$. For example, if $\ell=0$, then we find that the $X$-variety is (except some possible exceptional fibers) the graph of the map $f$. It also encodes the multivariable analogue of reflection across the sphere. The $X$-variety was defined by Forstnerič for the rationality result, and it has been extensively studied by D’Angelo, see [6, 7]. In fact, due to the connection of $k_{f}$ and the dimension of the fiber, some of the basic results about the $k_{f}$ integer follow directly from the work of D’Angelo, although we present direct proofs here. In particular D’Angelo developed a technique of quickly computing the fibers of the $X$-variety by associating to the map a certain matrix $C(w)$ whose nullspace gives the fiber of the $X$-variety at each point. Bounding the general hyperplane rank using Green’s hyperplane restriction theorem from commutative algebra was also used in a paper by Grundmeier, the second author, and Vivas [18] to prove a rigidity theorem for CR maps between hyperquadrics.
Let us then derive a few basic results about the $k_{f}$ integer. First we note that for any rational map, it is enough to check an open dense subset of hyperplanes, that is, the general hyperplane. A similar but more complicated statement holds for holomorphic maps defined in some domain, but we have no need for the complication. The space of affine hyperplanes in $\mathbb{C}^{n}$ has the topology of the Grassmannian (minus the hyperplane at infinity).